OpenStudy (anonymous):

When you're responding to a question, how do you make it draw nice-looking fractions?

OpenStudy (saifoo.khan):

\[\Huge \begin{array}l\color{red}{\text{C}}\color{orange}{\text{o}}\color{#9c9a2e}{\text{l}}\color{green}{\text{o}}\color{blue}{\text{r}}\color{purple}{\text{f}}\color{purple}{\text{u}}\color{red}{\text{l}}\color{orange}{\text{,}}\color{#9c9a2e}{\text{ }}\color{green}{\text{F}}\color{blue}{\text{r}}\color{purple}{\text{a}}\color{purple}{\text{c}}\color{red}{\text{t}}\color{orange}{\text{i}}\color{#9c9a2e}{\text{o}}\color{green}{\text{n}}\color{blue}{\text{s}}\color{purple}{\text{!}}\color{purple}{\text{!}}\color{red}{\text{!}}\color{orange}{\text{}}\end{array}\]

OpenStudy (anonymous):

No, I mean the horizontal line :)

OpenStudy (saifoo.khan):

\[\Huge \frac{Hi}{There}\]

OpenStudy (akshay_budhkar):



OpenStudy (saifoo.khan):

\[\Huge alax\frac{Like}{This?}\]

OpenStudy (anonymous):

\[Saifoo=\frac{likes}{rainbows}\]

OpenStudy (anonymous):

Yeah, like that. How do you do it?

OpenStudy (anonymous):

You must learn the way of the fractions

OpenStudy (anonymous):

Don't get too ahead of yourself

OpenStudy (anonymous):

alex you hv 3 options equation , draw and attach file the window shown here by akshay can be got by choosing the option draw am i right akshay

OpenStudy (akshay_budhkar):

yes you are :)

OpenStudy (anonymous):

Equation doesn't have a fraction button =]

OpenStudy (anonymous):

(x^2+3)/5

OpenStudy (anonymous):